What is the slope of the line: 5y-x=10
Answer by stanbon(75887)

What is the slope of the line: 5y-x=10
----
The slope is the amount y changes when x changes by 1.
----
5y = x + 10
---
y = (1/5)x + 2
----
Slope = 1/5
==================
Cheers,
Stan H.
==============